OpenStudy (anonymous):

The area of the parallelogram shown is twenty millimeters squared. What is the perimeter? Show your work.

OpenStudy (anonymous):

i will give a medal

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

dang this si a long answer

OpenStudy (esshotwired):

The area of a parallelogram is b which is the length of any base times a which is the corresponding altitude: \[b \times h=20mm ^{2}\] http://www.mathopenref.com/parallelogramarea.html First off, you can just take the squared part off the equation. Also, is there a picture to go with it?

OpenStudy (anonymous):

yah let me send it to you

OpenStudy (esshotwired):

It says I need a password and username :/

OpenStudy (anonymous):

never mind The area of the parallelogram shown is twenty millimeters squared. What is the perimeter? Show your work. A parallelogram is shown. The slanted sides are labeled thirteen millimeters. A dotted line extends from the lower right corner of the parallelogram to the right. This line intersects a second dotted line that stretches from the upper right corner of the parallelogram downward, intersecting the first dotted line to form a right angle. This dotted line is labeled ten millimeters.

OpenStudy (anonymous):

is this okay

OpenStudy (esshotwired):

Kinda. Let me draw it out and think about it

OpenStudy (anonymous):

i need this pronto

OpenStudy (esshotwired):

Ok so you know that 10 will be a in the equation so plug it in and solve that way for b. \[b \times 10=20mm ^{2}\] When you solve for b, b will equal 2. so all the sides of the shape equal 13+13+2+2=perimeter. So try solving the last part yourself

OpenStudy (anonymous):

what i don't get is the fact that the area is 20 and the perimeter is 10+13=23

OpenStudy (esshotwired):

10 is not part of the shape. It is just used to find the area so it is not counted as a side for the perimeter

OpenStudy (anonymous):

okay finish what you were doing then

OpenStudy (esshotwired):

I did, I solved for b which would be the base of the parallelogram. Then I gave you the equation to solve for the perimeter: 13+13+2+2=perimeter

OpenStudy (anonymous):

thank you so much!!!

OpenStudy (esshotwired):

your welcome! do you understand it better now?

OpenStudy (anonymous):

yah i think i will ace the question

OpenStudy (anonymous):

yout the bomb

OpenStudy (esshotwired):

lol thanks. hope u do well